CAVITY RING DOWN AT LOW TEMPERATURES: VIBRATIONAL OVERTONE ABSORPTION OF DEUTERATED METHANES AT 110 K

\underline{Y. PEREZ-DELGADO} and CARLOS E. MANZANARES, Department of Chemistry and Biochemistry, Baylor University, Waco, Texas, 76798.

The C-H stretch overtone absorptions of (Δυ = 5) of CH_{4}, CH_{3}D, CH_{2}D_{2}, and CHD_{3} have been obtained using the phase shift cavity ring down technique. A low temperature cell was designed and adapted to the CRD experiment to measure absorption bands at any temperature between 10 K and 298 K. The spectra were obtained at room temperature and at 110 K. The partially resolved rotational spectrum of CHD_{3} that included the transitions 5ν_{1} and 4ν_{1} + 2ν_{5} was obtained and compared with the calculated spectrum. The unresolved vibrational bands around the Δυ = 5 region of CH_{2}D_{2}, CH_{3}D, and CH_{4} were also obtained at several temperatures between 100K and 298 K. The integrated absorption was calculated as a function of the density of the gas samples and used to obtain the cross section and the oscillator strength of the transition for each molecule.